AGENTJSVSL:

I just need someone to check that I got the correct answer. I chose B. The average IQ score follows a Normal distribution, with a mean of μ = 100 and a standard deviation of σ = 15. What is the probability that the mean IQ score of 25 randomly selected people will be greater than 105? 0 0.0478 0.50 0.68 0.9522 Your input is highly appreciated. Thanks in advance.

Shadow:

You have to find the z-score for 105 then obtain the tail probability since it is less than half of the distribution from here: http://math-stat.net/unit%20normal%20table.pdf

AGENTJSVSL:

Oh I calculated using a normal Pieulative probability..should I show my work so that you can check?

AGENTJSVSL:

Marsulative*

AGENTJSVSL:

I think the filter won't let me say Cu mulative

Shadow:

Yes it won't.

AGENTJSVSL:

Should I show my work so that you can check?

Shadow:

Sure

AGENTJSVSL:

I did lower limit --> 105 upper limit --> 999, 999 sample mean --> 100 standard dev of sample --> σ/√n = 3 normalcdf(105, 999999, 100, 3) =0.0477

Shadow:

Hmm, the way I usually do these is as follows \[\sigma_{m} = \frac{ \sigma }{ \sqrt n} = \frac{ 15 }{ \sqrt 25 } = 3\] \[z = \frac{ M - \mu }{ \sigma_{m} } = \frac{ 105-100 }{ 3 } = \frac{ 5 }{ 3 } = 1.66\] Going along with the pdf I showed you, the tail probability in accordance with the z-score of 1.66 is 0.0485

AGENTJSVSL:

Smart approach that works as well. Thank you for taking the time to help me out!